# Turbulence-induced anti-Stokes flow: experiments and theory

### v1

@@ 4.4 A model based on RDT
The previous calculations only apply to the turbulence far away from the air-water interface (but affected by the Stokes drift, because of the scale separation $L\ll\lambda$). In order to take into account the blocking effect of the air-water interface, where we assume that the interface affects the turbulence essentially as a frictionless wall for depths of $O(L)$ (because the air-water interface has a large density contrast), this effect can be accounted for by adding an irrotational correction to the homogeneous turbulent velocity field, as done before by Hunt & Graham (1978) and Teixeira & Belcher (2002). Note that if the waves that generate the Stokes drift are irrotational (which is true to a good degree of approximation), then this irrotational correction remains irrotational. Given this, the turbulent velocity components affected both by the Stokes drift and by blocking can be expressed as 2D Fourier integrals along the horizontal directions,
@@
$\overline{\hat{u}_{i0}^{(H)}({\boldsymbol{\kappa}}_{0})\hat{u}_{j0}^{(H)}({\boldsymbol{\kappa}}_{0}^{\prime})}=\Phi_{ij}^{(H)}({\boldsymbol{\kappa}}_{0})\delta({\boldsymbol{\kappa}}_{0}-{\boldsymbol{\kappa}}_{0}^{\prime}),$ (28)
where ${\boldsymbol{\kappa}}_{0}=(\kappa_{10},\kappa_{20},\kappa_{30})$, $\Phi_{ij}^{(H)}({\boldsymbol{\kappa}}_{0})$ is the three-dimensional spectrum of the initial homogeneous and isotropic turbulence, and $\delta$ is the three-dimensional Dirac delta, to show that $\overline{u^{\prime}_{1}u^{\prime}_{3}}$ is given by
@@
where $q$ is a characteristic RMS turbulent velocity and $L$ is the longitudinal integral length scale. The constants $g_{1}\approx 0.558$ and $g_{2}\approx 1.196$ ensure that the definitions of $q$ and $L$ are consistent.
Note that while a number of simplifying assumptions have been made which mean that full quantitative agreement with observations is not to be expected, the above theory involves no fitting or particular tailoring to our specific system. The turbulence is described parameterised by just two scalars pertaining to the homogeneous bulk, $q$ and $L$, a gross simplification which nevertheless captures the most essential features. Importantly, these are quantities obtainable from point measurements in experiments as well as in the field.
@@ 4.5 Input parameters
where the TKE $e$ is found in Table 3.
@@
The displacement $\beta$ defined in (21) is determined by the duration and extent of interaction between waves and turbulence in the time between generation by the active grid, and measurement a distance $L_{\mathrm{FOV}}$ downstream. For a Gaussian wave packet in cases 1.A1.D, van den Bremer et al. (2019) showed that the final value of $\beta$ after the passage of the group, $\beta_{\mathrm{f}}$ (subscript ‘$\mathrm{f}$’: final), takes the form
$\beta_{\mathrm{f}}(z)=4\sqrt{\pi}\sigma_{0}k_{0}\epsilon_{0\mathrm{p}}^{2}\mathrm{e}^{2k_{0}z},$ (34)
where $\sigma_{0}$ is the spatial standard deviation of the Gaussian wave group and $\epsilon_{0\mathrm{p}}$ the peak steepness. By definition of wave group, $\sigma_{0}=c_{g}(k_{0})\tau_{0}$ (see (2) and Table 3). The expression presumes that the turbulence moving downstream from the active grid has interacted with the whole wave group before reaching the field of view; this is true in our Experiment 1.
@@
$\beta=t_{\mathrm{travel}}\frac{\mathrm{d}u_{s}}{\mathrm{d}z}=\frac{2L_{\mathrm{FOV}}}{U_{0}}k_{0}\epsilon_{0}^{2}c(k_{0})\mathrm{e}^{2k_{0}z}.$ (35)
@@
4.6 RDT results
Figure 7: Results from RDT: (a) profile of the normalized shear stress $\overline{u^{\prime}w^{\prime}}/q^{2}$ as a function of depth $k_{0}z$; (b) profile of the normalized vertical derivative of the shear stress; (c) profile of the normalized shear stress as a function of $\beta$ for $k_{0}z=-0.4335$, the depth where $\overline{u^{\prime}w^{\prime}}/q^{2}$ attains its maximum magnitude. The dashed line is the 1:1 line, illustrating a linear dependence. (d) RDT estimates of the anti-Stokes velocity profile after the passage of the wave groups, corresponding to figure 3(b).
@@ 4.6 RDT results
Next, figure 7(c) shows the dependence of the shear stress at its maximum on $\beta$ (which is proportional to the square of the wave slope $\epsilon_{0}$). The vertical derivative of the shear stress at the surface, to which the current at the surface must be proportional, is clearly proportional to this maximum. This implies that the current at the surface (departing from a quiescent flow) must be proportional to $\epsilon_{0}^{2}$, which is consistent with the results of figure 5 for cases 2.A and 2.B.
@@
While the primary goal of our RDT model is to elucidate the physical process behind our observations, we have also used (29) to obtain a rough estimate of the current resulting from the wave groups in Experiment 1 (Cases 1.A1.D) encountering the incoming turbulence. In order to obtain quantitative values, we use a simplistic procedure: noting that $\partial_{z}\overline{u^{\prime}w^{\prime}}\propto\epsilon^{2}$, we estimate the current after the passage of the group as
@@
$\approx-(\partial_{z}\overline{u^{\prime}w^{\prime}})_{\text{RDT}}T_{\mathrm{int}}$ (36)
inserting the calculated Reynolds stress vertical gradient shown in figure 7(b) multiplied by $k_{0}q^{2}$, using values for $k_{0},q=2e/3$ and $L=L_{x}^{x}$ from table 3 and the “effective interaction time” from (6). The result is shown in figure 7(d).
There is little reason to expect close agreement between our measured $\Delta U(z)$ in Experiment 1 and $U_{\text{RDT}}(z)$ because the assumptions behind (36) can only hold very early in the ‘spin-up’ period; the relation assumes that $\overline{u^{\prime}w^{\prime}}$ retains the vertical shape shown in figure 7(a) throughout the duration of the interaction between wave group and turbulence. Gradually, however, the turbulent flow field begins to ‘push back’ via the growth of other terms in (12) until $\partial_{z}\overline{u^{\prime}w^{\prime}}$ eventually becomes very small. We are unfortunately unable to study this transient process in detail in our present experiment — it is a highly relevant question for future study, experimentally and numerically (e.g. in the vein of Guo & Shen, 2013).
Despite the naïveté of (36) and the numerous assumptions made when applying RDT, the agreement with the measurements of $\Delta U$ in figure 3(b) is not only qualitatively but also quantitatively reasonable. The predictions vary with depth notably faster than the measured currents, which we conjecture is a consequence of the smoothing effect of turbulence as it develops beyond the linear-growth regime described by (36). Agreement is comparatively poor in Case 1.B where $U_{\text{RDT}}$ is far smaller than the corresponding $\Delta U$, which we cannot explain at present (Note that this same case displays a surprisingly strong modification of the underlying turbulence considering that the TKE of Case 1.B is intermediate, as seen in figure 4 of Smeltzer et al. (2023)).
The key conclusion to be drawn is perhaps that the change in Eulerian current can indeed be ascribed to the interaction between waves and turbulence, to wit, the gradual tilting and stretching of vortices by the Stokes drift as previously studied by Teixeira & Belcher (2002).
@@ 5 Conclusions
We have presented experimental evidence that an Eulerian-mean flow directed opposite to the waves’ propagation direction is created when waves propagate atop a turbulent flow, and argued via two different theoretical approaches how the current is the result of waves and turbulence interacting.
@@
We present two separate theoretical models which can describe the quasi-steady state and the transition period, respectively. For the latter situation an approximate relation between the mean current shear $\mathrm{d}{\boldsymbol{u}}/\mathrm{d}z$ and the Stokes drift gradient $\mathrm{d}u_{s}/\mathrm{d}z$ is derived following Pearson (2018), valid nearest the surface where $u_{s}$ is significant. The relation involves the turbulent variances $\overline{u^{\prime}u^{\prime}}$ and $\overline{u^{\prime}w^{\prime}}$ and shows good enough agreement with experiments to inspire confidence in its use, in the cases where underlying assumptions are satisfied.
A further approximate relation is adapted from Pearson (2018) and relates the rate of change of the Eulerian-mean current to the Reynolds stress $\partial_{z}\overline{u^{\prime}w^{\prime}}$. Hypothesising that the underlying mechanism is the action of Stokes drift on turbulent eddies, Rapid Distortion Theory (RDT) is used to estimate $\overline{u^{\prime}w^{\prime}}(z)$ and hence the time-varying ${\boldsymbol{u}}(z,t)$ in the early phase of interaction. The predictions of RDT are compared with the measurements of ${\boldsymbol{u}}(z)$ due to passing wave groups, by using measured values of turbulent kinetic energy, integral length scale, and group envelope width. Qualitative and quantitative agreement with measurements is found, despite a series of simplifying assumptions.
@@ Acknowledgements
We have benefited from discussions with a number of colleagues, particularly Victor Shrira and Nick Pizzo. We thank Marc Buckley for invaluable input and for sharing data-analysis code, and Masoud Asadi, Leon Li and Pim Bullee for assistance with the experiments.
@@
The work was co-funded by the Research Council of Norway (iMOD, 325114, and FRIPRO, WallMix, 288046) and the European Union (ERC StG, GLITR, 101041000 and ERC CoG, WaTurSheD, 101045299). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@
Primary contributions are as follows. S.Å.E.: Principal writer, theory, project co-ordination, supervision. O.R.: Planned and performed experiment 3, conducted the majority of the data analysis. B.K.S.: Planned and performed experiments 1 and 2, first discovered the change in current in the experimental data. M.A.C.T.: Developed the Rapid Distortion Theory. T.S.vdB.: Theory development, discussions. K.S.M.: Error and convergence analysis. R.J.H.: Experimental and theoretical contributions, supervision and discussions. S.Å.E. and O.R. should be considered to have made equal contributions. All authors contributed significantly to writing the manuscript.
@@ Appendix A Active grid settings
The settings used for the active grid in the different cases 1.A3.B are listed in table 4. The active grid contains two sets of bars; vertical and horizontal. The phrase static, open indicates that the grid bars are not moving but in a fully open position, causing as little blockage of the inflow as possible. Cases labelled random rotation means that grid bars are actuated randomly with a top-hat distribution centred at $\overline{f_{G}}$ spread over ±$\pm\delta f_{G}$. The rotation direction is also random with an equal likelihood of clockwise and counter-clockwise rotation. For case 3.B the grid bars are flapped ±$\pm 60^{\circ}$ about the fully open position. Here, a flap motion occurs at random intervals uniformly distributed between $5$ and $10$ seconds. Each bar is flapped independently.
@@
1.A 00 static, open static, open
@@
3.B - static, open random flapping ±$\pm 60^{\circ}$

### v2

@@ 4.4 A model based on RDT
The previous calculations only apply to the turbulence far away from the air-water interface, at least a distance $\sim L$ where $L$ is a characteristic integral length scale (but affected by the Stokes drift, because of the scale separation $L\ll\lambda$). In order to take into account the blocking effect of the air-water interface, where we assume that the interface affects the turbulence essentially as a frictionless wall for depths of $O(L)$ (because the air-water interface has a large density contrast), this effect can be accounted for by adding an irrotational correction to the homogeneous turbulent velocity field, as done before by Hunt & Graham (1978) and Teixeira & Belcher (2002). Note that if the waves that generate the Stokes drift are irrotational (which is true to a good degree of approximation), then this irrotational correction remains irrotational. Hence, the turbulent velocity components affected both by the Stokes drift and by blocking can be expressed as two-dimensional Fourier integrals along the horizontal directions,
@@
$\overline{\hat{u}_{i0}^{(H)}({\boldsymbol{\kappa}}_{0})^{*}\hat{u}_{j0}^{(H)}({\boldsymbol{\kappa}}_{0}^{\prime})}=\Phi_{ij}^{(H)}({\boldsymbol{\kappa}}_{0})\delta({\boldsymbol{\kappa}}_{0}-{\boldsymbol{\kappa}}_{0}^{\prime}),$ (28)
where an asterisk denotes the complex conjugate, $\Phi_{ij}^{(H)}({\boldsymbol{\kappa}}_{0})$ is the three-dimensional spectrum of the initial homogeneous and isotropic turbulence and $\delta$ is the three-dimensional Dirac delta, to show that $\overline{u^{\prime}_{1}u^{\prime}_{3}}$ is given by
@@
where $q$ is a characteristic r.m.s. value of a turbulent velocity component and $L$ is the longitudinal integral length scale. The constants $g_{1}\approx 0.558$ and $g_{2}\approx 1.196$ ensure that the definitions of $q$ and $L$ are consistent.
Note that while a number of simplifying assumptions have been made which mean that full quantitative agreement with observations is not to be expected, the above theory involves no fitting or particular tailoring to our specific system. The turbulence is described by just two scalar parameters pertaining to the homogeneous bulk, $q$ and $L$, a crude simplification which nevertheless is expected to capture the most essential features. Importantly, these are quantities obtainable from point measurements in experiments as well as in the field.
@@ 4.5 Input parameters
assuming isotropic turbulence, where the TKE $e$ is found in Table 2.
@@
The displacement $\beta$ defined in (21) is determined by the duration and extent of interaction between waves and turbulence in the time between generation by the active grid, and measurement a distance $L_{\mathrm{FOV}}$ downstream. For a Gaussian wave packet like those in cases 1.A1.D, van den Bremer et al. (2019) showed that the final value of $\beta$ after the passage of the group, $\beta_{\mathrm{f}}$ (subscript $\mathrm{f}$’: final), takes the form
$\beta_{\mathrm{f}}(z)=4\sqrt{\pi}\sigma_{0}k_{0}(k_{0}a_{\mathrm{p}})^{2}\mathrm{e}^{2k_{0}z},$ (34)
where $\sigma_{0}$ is the spatial standard deviation of the Gaussian wave group and $k_{0}a_{\mathrm{p}}$ the peak steepness. By definition of wave group, $\sigma_{0}=c_{g}(k_{0})\tau_{0}$ (see (2) and Table 2). The expression presumes that the turbulence moving downstream from the active grid has interacted with the whole wave group before reaching the field of view; this is true in our Experiment 1.
@@
$\beta=t_{\mathrm{travel}}\frac{\mathrm{d}u_{s}}{\mathrm{d}z}=\frac{2k_{0}L_{\mathrm{FOV}}}{U_{0}}(ak_{0})^{2}c(k_{0})\mathrm{e}^{2k_{0}z}.$ (35)
@@
4.6 The RDT results
Case $T_{\mathrm{int}}$ $T_{\mathrm{int}}/T_{0}$ $\beta_{\mathrm{f}}(0)$
$(\mathrm{s})$
1.A $4.3$ $6.6$ $3.3$
1.B $4.6$ $7.0$ $3.5$
1.C $5.1$ $7.6$ $2.2,4.7$
1.D $5.0$ $7.6$ $4.6$
2.A.1 $29.2$ $35.7$ $3.6,8.6,14.2$
2.A.2 $29.2$ $50.6$ $13.9,22.3,27.2$
2.B.1 $29.2$ $35.9$ $2.1,6.3,12.7$
2.B.2 $29.2$ $50.4$ $12.1,17.8,27.1$
3.A $46.1$ $82.5$ $12.2,32.6$
3.B $46.1$ $82.8$ $12.2,32.7$
Table 4: Derived wave quantities for use in RDT analysis. For cases 1.A1.D, $T_{\mathrm{int}}=\sqrt{\pi}\tau_{0}$ is used with $\tau_{0}$ from Eq.0(5), while for cases 2.A3.B, $T_{\mathrm{int}}=L_{\mathrm{FOV}}/U_{0}$; $\beta_{\mathrm{f}}(0)$ is found from Eqs. (34) and (35) for groups and regular waves, respectively. Note that it is related to $T_{\mathrm{int}}$ via Eq.0(39).
Figure 7: Results from RDT: (a) profile of the normalised shear stress $\overline{u^{\prime}w^{\prime}}/q^{2}$ as a function of depth $k_{0}z$; (b) profile of the normalized vertical derivative of the shear stress; (c) the normalised shear stress as a function of $\beta$ for $k_{0}z=-0.4335$, the depth where $\overline{u^{\prime}w^{\prime}}/q^{2}$ attains its maximum magnitude. The dashed line is the 1:1 line, illustrating a linear dependence. (d) The RDT estimates of the anti-Stokes velocity profile after the passage of the wave groups, corresponding to figure 3(b).
@@ 4.6 The RDT results
Next, figure 7(c) shows the dependence of the shear stress at its maximum on $\beta$ (which is proportional to the square of the wave slope $ak_{0}$ as Eqs. (34) and (35) show). The vertical derivative of the shear stress at the surface, to which the current at the surface must be proportional, is clearly proportional to this maximum. This implies that the current at the surface (departing from a quiescent flow) must be proportional to $(ak_{0})^{2}$, which is consistent with the results of figure 5 for cases 2.A and 2.B.
@@
4.6.1 Effective interaction time
While the primary goal of our RDT model is to elucidate the physical process behind our observations, we have also used (29) to obtain a rough estimate of the current resulting from the wave groups in Experiment 1 (Cases 1.A1.D) encountering the incoming turbulence. In order to obtain quantitative values, we use a simplified procedure: noting that $\partial_{z}\overline{u^{\prime}w^{\prime}}\propto(ak_{0})^{2}$, we estimate the current after the passage of the group as
@@ 4.6.1 Effective interaction time
$\approx-(\partial_{z}\overline{u^{\prime}w^{\prime}})_{\text{RDT}}T_{\mathrm{int,grp}}$ (36)
where we have introduced an “effective interaction time” for a wave group,
$T_{\mathrm{int,grp}}=\frac{1}{a_{\mathrm{p}}^{2}}\int_{-\infty}^{\infty}\tilde{a}(t)^{2}\mathrm{d}t,$ (37)
an estimate for the duration for which the turbulent current has been influenced by the waves before it reaches the point where it is measured, taking into account that the wave amplitude varies throughout the group. As seen by the turbulence, $\tilde{a}(t)$ is a Gaussian with $\tau_{0}$ from Eq. (5) replacing $\tau$ in (4), inserting which into (37) gives
$T_{\mathrm{int,grp}}=\sqrt{\pi}\tau_{0}.$ (38)
We next insert into (36) the calculated Reynolds stress vertical gradient shown in figure 7(b) (multiplied by $k_{0}q^{2}$ to obtain a dimensional value), using values for $k_{0},q=2e/3,L=L_{x}^{x}$ and $T_{\mathrm{int,grp}}$ from table 4. The resulting estimated velocity profiles are shown in figure 7(d).
It is worth noting the relation between the effective interaction time and the fluid displacement $\beta$. For regular waves, the turbulence interacts constantly with waves throughout its passage from grid to field of view, so the interaction time in the presence of regular waves is $T_{\mathrm{int,reg}}=L_{\mathrm{FOV}}/U_{0}$. Using Eq. (34) and noting that for deep-water waves, $\sigma_{0}=c_{g}\tau_{0}={\textstyle\frac{1}{2}}c\tau_{0}$, we find that for regular waves as well as wave groups it holds that
$\beta(z)=2k_{0}T_{\mathrm{int}}u_{s}(z).$ (39)
4.6.2 Comparison with experiment
There is little reason to expect close quantitative agreement between our measured $\Delta U(z)$ in Experiment 1 and $U_{\text{RDT}}(z)$ because the assumptions behind (36) can only hold very early in the ‘spin-up’ period; the relation assumes that $\overline{u^{\prime}w^{\prime}}$ retains the vertical shape shown in figure 7(a) throughout the duration of the interaction between wave group and turbulence. Gradually, however, the turbulent flow field begins to ‘push back’ via the growth of other terms in (12) until $\partial_{z}\overline{u^{\prime}w^{\prime}}$ eventually becomes very small. Investigating this transient process in detail is a highly relevant question for future study, experimentally and numerically (e.g. in the vein of Guo & Shen, 2013).
Despite the naïveté of (36) and the numerous assumptions made when applying RDT, the agreement with the measurements of $\Delta U$ in figure 3(b) is not only qualitatively but also quantitatively reasonable. The predictions vary with depth notably faster than the measured currents, which we conjecture is a consequence of the smoothing effect of turbulence as it develops beyond the linear-growth regime described by (36), a turbulent mixing which RDT by construction cannot account for. Agreement is comparatively poor in Case 1.B where $U_{\text{RDT}}$ is far smaller than the corresponding $\Delta U$, which we cannot explain at present (note that this same case displays a surprisingly strong modification of the underlying turbulence considering that the TKE of Case 1.B is intermediate, as seen in figure 4 of Smeltzer et al. (2023)).
The key conclusion to be drawn is perhaps that the change in Eulerian current can indeed be ascribed to the interaction between waves and turbulence, to wit, the gradual tilting and stretching of vortices by the Stokes drift, as previously studied by Teixeira & Belcher (2002).
@@ 5 Conclusions
We have presented experimental evidence that an Eulerian-mean flow directed opposite to the waves’ propagation direction is created near the water surface when waves propagate atop a turbulent flow, and argued via two different theoretical approaches how the current is the result of waves and turbulence interacting.
@@
We present two separate theoretical models which can describe the quasi-steady state and the transition period, respectively. For the latter situation an approximate relation between the mean current shear $\mathrm{d}{\boldsymbol{u}}/\mathrm{d}z$ and the Stokes drift gradient $\mathrm{d}u_{s}/\mathrm{d}z$ is derived following Pearson (2018), valid nearest the surface where $u_{s}$ is significant. The relation involves the turbulent variances $\overline{u^{\prime}u^{\prime}}$ and $\overline{w^{\prime}w^{\prime}}$ and shows good enough agreement with experiments to inspire confidence in its use, in the cases where underlying assumptions are satisfied.
A further approximate relation is adapted from Pearson (2018) and relates the rate of change of the Eulerian-mean current to the Reynolds stress $\partial_{z}\overline{u^{\prime}w^{\prime}}$. Hypothesising that the underlying mechanism is the action of Stokes drift on turbulent eddies, Rapid Distortion Theory (RDT) is used to estimate $\overline{u^{\prime}w^{\prime}}(z)$ and hence the time-varying ${\boldsymbol{u}}(z,t)$ in the early phase of interaction. The predictions of RDT are compared with the measurements of ${\boldsymbol{u}}(z)$ due to passing wave groups, by using measured values of TKE, integral length scale, and group envelope width. Qualitative and quantitative agreement with measurements is found, despite a series of simplifying assumptions.
@@ Acknowledgements
We have benefited from discussions with a number of colleagues, particularly Victor Shrira and Nick Pizzo. We thank Marc Buckley for invaluable input and for sharing data-analysis code, and Masoud Asadi, Leon Li and Pim Bullee for assistance with the experiments. We are grateful to Rafael Bölsterli for feedback on the manuscript.
@@
The work was co-funded by the Research Council of Norway (iMOD, 325114) and the European Union (ERC StG, GLITR, 101041000 and ERC CoG, WaTurSheD, 101045299). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. M.A.C.T. was supported by Portuguese national funds through FCT/MCTES (PIDDAC), for the ResearchUnits CEFT:UID/00532, and ALiCE: LA/P/0045/2020 (doi: 10.54499/LA/P/0045/2020).
@@
Primary contributions are as follows. S.Å.E.: Principal writer, theory, project co-ordination, supervision. O.R.: Planned and performed experiment 3, conducted the majority of the data analysis. B.K.S.: Planned and performed experiments 1 and 2, first discovered the change in current in the experimental data. M.A.C.T.: Developed the Rapid Distortion Theory. T.S.vdB.: Theory development, discussions. K.S.M.: Error and convergence analysis. R.J.H.: Supervision of experiments, turbulence-theoretic contributions and discussion. S.Å.E. and O.R. should be considered to have made equal contributions. All authors contributed significantly to the writing of the manuscript.
Data availability statement
The experimental data from experiment 1 are available at doi:10.18710/R0I0RW, and for experiments 2 and 3 from doi:10.18710/ZZPUNJ.
@@ Appendix A Active grid settings
The settings used for the active grid in the different cases 1.A3.B are listed in table 5. The active grid contains two sets of bars; vertical and horizontal. The phrase static, open indicates that the grid bars are not moving but in a fully open position, causing as little blockage of the inflow as possible. Cases labelled random rotation means that grid bars are actuated randomly with a top-hat distribution centred at $\overline{f_{G}}$ spread over $\pm\delta f_{G}$. The rotation direction is also random with an equal likelihood of clockwise and counter-clockwise rotation. For case 3.B the grid bars are flapped $\pm 60^{\circ}$ about the fully open position. Here, a flap motion occurs at random intervals uniformly distributed between $0.5$ and $1.0$ seconds. Each bar is flapped independently.
@@
1.A 0 static, open static, open
@@
3.B - static, open random flapping $\pm 60^{\circ}$